OpenStudy (anonymous):

Give me the answer and right the steps (complete the table of values) 2x - 5y =-10

OpenStudy (anonymous):

What is the table of values? Are you uploading an image or adding additional text for reference?

OpenStudy (anonymous):

X y 2 0 4

OpenStudy (anonymous):

just plug the y coordinate to get the x coord. so you would plug in 2 to get the x, then plug in 0 again to get that x, and so on. that's how you complete it

OpenStudy (anonymous):

I dont know math at all that why I came here so u can solve it

OpenStudy (anonymous):

sure. so you use the equation 2x - 5y =-10. you have three values on your y chart: 2, 0 and 4. let's work with 2 first ok? so 2x-5y=-10 x-2y=-10 <-plug in 2 for y, because you have y coordinate x-2(2)=-10 x-4=-10 x=-10+4 x=-6. so the first pair of coordinates are (-6,2). do you need me to do the rest?

OpenStudy (anonymous):

I really only need you to write the answers im doing a test